### v2

# Low-Rank Tensor Decomposition over Finite Fields

Abstract
We show that finding rank-1, rank-2, and rank-3 decompositions of a 3D tensor over a fixed finite field can be done in polynomial time. However, if some cells in the tensor are allowed to have arbitrary values, then rank-2 is NP-hard over the integers modulo 2. We also explore rank-1 decomposition of a 3D tensor and of a matrix where some cells are allowed to have arbitrary values.
1 Introduction
Given a $p\times q\times s$ tensor (multidimensional array) $T$, a rank-$R$ decomposition is a collection of variables $A_{a,b},B_{c,d},C_{e,f}$ such that
$T_{i,j,k}=\sum_{r=0}^{R-1}A_{r,i}B_{r,j}C_{r,k}\ \forall i,j,k.$
This equation is also written as $T=\sum_{r=0}^{R-1}A_{r,:}\times B_{r,:}\times C_{r,:}$, where $M_{r,:}$ denotes the slice $[M_{r,j}]_{j}$ and $\times$ denotes the outer product: for arbitrary tensors $X$ and $Y$, $(X\times Y)_{i_{0},\dots i_{k-1},j_{0},\dots j_{\ell-1}}:=X_{i_{0},\dots i_{k-1}}Y_{j_{0},\dots j_{\ell-1}}$.
Tensor decomposition is at the heart of fast matrix multiplication, a problem that is the primary bottleneck of numerous linear algebra and graph combinatorics algorithms, such as matrix inversion and triangle detection. All asymptotically fast algorithms for matrix multiplication use a divide-and-conquer scheme, and finding an efficient scheme is equivalent to finding a decomposition of a certain tensor with low rank [2].
While significant progress has been made on this problem over the decades, most of the asymptotically fastest algorithms discovered use indirect constructions and convergence arguments that result in enormously large divide-and-conquer schemes, making the constant factors for these algorithms far too big for them to be useful in practice. For this reason, there has also been interest in finding explicit decompositions for the tensors of small divide-and-conquer schemes. With the introduction of AlphaTensor [3] and flip graphs [5] just over a year ago, several new low-rank decompositions have been discovered, especially over finite fields of characteristic 2 [1].
Decompositions using only integers are especially preferred, since they avoid having to use floating point arithmetic, and most decompositions of small divide-and-conquer schemes that have been found have happened to use only integers. This makes it useful to search for decompositions over finite fields (e.g., integers modulo 2), as doing so makes the search space finite and allows one to rule out large classes of integer solutions.
It is known that deciding whether a given tensor has a decomposition of rank $\leq R$ is NP-hard over finite fields via a reduction from 3SAT [4]. However, since we only want low-rank decompositions, we are more interested in hardness results for small fixed $R$.
1.1 Results
Our main result in this report is a polynomial-time algorithm for 3D tensor decomposition with rank $\leq 3$. For simplicity of asymptotic analysis, we focus on $n\times n\times n$ tensors $T$, although our algorithm easily generalizes to non-cube tensors.
Theorem 1.
For $R\leq 3$, finding a rank-$R$ decomposition over a finite field $\mathbb{F}$ of an $n\times n\times n$ tensor, or determining that it does not exist, can be done in $O(f(|\mathbb{F}|,R)n^{3})$ time, for $f$ polynomial in $|\mathbb{F}|$ and exponential in $R$.
However, when allowing certain cells of the target tensor to have arbitrary value (“wildcards”), we have a hardness result over the integers mod 2 via a reduction from Not-All-Equal 3SAT (NAE-3SAT):
Theorem 2.
If some cells of a $p\times q\times s$ tensor $T$ are allowed to have arbitrary value, then determining if $T$ has a rank-2 decomposition over $\mathbb{Z}/2\mathbb{Z}$ is NP-hard.
We also sketch out polynomial-time algorithms for rank-1 decomposition with wildcards over a 3D tensor and a matrix, over an arbitrary fixed finite field.
2 Without wildcards
In this section we describe our algorithm for Theorem 1. The following observations of rank-$\leq 2$ decomposition over $\mathbb{Z}/2\mathbb{Z}$ serve as inspiration:
  • Rank-1 is easy since $abc=1$ forces $a=1$, $b=1$, and $c=1$, so $T=a\times b\times c$ if and only if $(a_{i}=\lor_{j,k}T_{i,j,k},\ b_{j}=\lor_{i,k}T_{i,j,k},\ c_{k}=\lor_{i,j}T_{i,j,k})$ is a solution.
  • Rank-2 at first glance is not so clear, but we notice that if $T=a_{0}\times b_{0}\times c_{0}+a_{1}\times b_{1}\times c_{1}$, then all slices $T_{i,:,:}$ are of the form $u(b_{0}\times c_{0})+v(b_{1}\times c_{1})$ for $u,v\in\mathbb{Z}/2\mathbb{Z}$, so each $T_{i,:,:}$ can only be one of four possibilities. These slices also have rank at most 2, so one could maybe do exhaustive casework over the ranks of these slices to find a rank-2 decomposition of $T$ (or prove that one does not exist).
Our algorithm generalizes to rank-3 and arbitrary finite fields. Suppose we want to find a rank-$R$ decomposition of the $n\times n\times n$ tensor $T$ over the finite field $\mathbb{F}$, i.e., $T_{i,j,k}=\sum_{r=0}^{R-1}A_{r,i}B_{r,j}C_{r,k}\ \forall i,j,k$. Our main insight is to fix every possible $A$ and then solve for $B,C$ in a clever manner. To ensure we do not enumerate too many $A$ (specifically, the number of possible $A$ we search over is $O(1)$ w.r.t. $n$), we extract a basis on the slices $T_{i,:,:}$.
Flatten each $T_{i,:,:}$ to form the $n\times n^{2}$ matrix $\overline{T}:=\left[T_{i,:,:}\texttt{.flatten}()\right]_{i}$111 We use numpy.ndarray.flatten to flatten an $n\times n$ matrix into a $n^{2}$-length vector, with elements in row-major order. For example, $\begin{bmatrix}a&b\\ c&d\end{bmatrix}\texttt{.flatten}()=\begin{bmatrix}a&b&c&d\end{bmatrix}$. ; run Gaussian elimination to get $\overline{T}=M\mathrm{rref}\left(\overline{T}\right)$, where rref denotes reduced row-echelon form and $M$ is invertible; let $J$ be the number of nonzero rows of $\mathrm{rref}\left(\overline{T}\right)$, which is equal to $\mathrm{rank}\left(\overline{T}\right)$. Then we can collect basis slices to form the tensor $T^{\prime}=\left[\mathrm{rref}\left(\overline{T}\right)_{j,:}\texttt{.reshape}(n,n)\right]_{0\leq j<J}$222 We use numpy.ndarray.reshape to reshape a $n^{2}$-length vector into a $n\times n$ matrix, with elements in row-major order. For example, $\begin{bmatrix}a&b&c&d\end{bmatrix}\texttt{.reshape}(2,2)=\begin{bmatrix}a&b\\ c&d\end{bmatrix}$. , yielding $T_{i,:,:}=\sum_{0\leq j<J}M_{i,j}T^{\prime}_{j,:,:}$ and $T^{\prime}_{j,:,:}=\sum_{0\leq i<n}(M^{-1})_{j,i}T_{i,:,:}$. These relations allow us to convert between decompositions of $T$ and $T^{\prime}$:
$T^{\prime}=\sum_{r=0}^{R-1}U_{r}\times B_{r,:}\times C_{r,:}$ $\Rightarrow T=\sum_{r=0}^{R-1}\left[\sum_{0\leq j<J}M_{i,j}U_{r,j}\right]_{i}\times B_{r,:}\times C_{r,:}$
$T=\sum_{r=0}^{R-1}A_{r}\times B_{r,:}\times C_{r,:}$ $\Rightarrow T^{\prime}=\sum_{r=0}^{R-1}\left[\sum_{0\leq i<n}(M^{-1})_{j,i}A_{r,i}\right]_{i}\times B_{r,:}\times C_{r,:}$
Thus, finding a rank-$R$ decomposition of $T$ is equivalent to finding a rank-$R$ decomposition of $T^{\prime}$. We can then rewrite $T^{\prime}=\sum_{r}U_{r}\times B_{r}\times C_{r}$ into the block matrix form $\begin{bmatrix}T^{\prime}_{j,:,:}\end{bmatrix}_{0\leq j<J}=U^{\intercal}\begin{bmatrix}B_{r,:}\times C_{r,:}\end{bmatrix}_{r}$.
The upshot of constructing $T^{\prime}$ is that if there exists $A,B,C$ s.t. $T=\sum_{r}A_{r}\times B_{r,:}\times C_{r,:}$, then $\forall i,\ T_{i,:,:}\in\textrm{span}\left\{B_{r,:}\times C_{r,:}\right\}_{0\leq r<R}$, so $\left\{T_{i,:,:}\right\}_{0\leq i<n}$ spans a linear subspace of $\mathbb{F}^{n\times n}$ with rank at most $R$. This forces $T^{\prime}$ to have at most $R$ many slices, i.e., $J\leq R$, so there are at most $|\mathbb{F}|^{R^{2}}$ possible $U$ to fix. Thus, we avoid exponential dependence on $n$.
If $J=0$, then $T$ must be all-zero, so it trivially has a rank-$R$ decomposition (e.g., with all terms set to 0). Otherwise, for every possible $U$, run Gaussian elimination on $\Xi:=\left[\begin{array}[]{c|c}U^{\intercal}&\begin{bmatrix}\vdots\\ T^{\prime}_{j,:,:}\\ \vdots\end{bmatrix}_{0\leq j<J}\end{array}\right]$ to obtain
$\Xi^{\prime}:=\left[\begin{array}[]{c|c}\Omega&\begin{bmatrix}\vdots\\ D_{j}\\ \vdots\end{bmatrix}_{0\leq j<J}\end{array}\right]$, for $\Omega:=\mathrm{rref}\left(U^{\intercal}\right)$ and some $n\times n$ matrices $D_{j}$.
If there is some row $r$ s.t. $\Omega_{r,:}$ is all-zeros but $D_{r}$ is not, then there is no solution. Otherwise, since all all-zero rows of a rref matrix must be at the bottom, we can remove everything except the first $N$ rows of $\Xi^{\prime}$, where $N$ is the number of nonzero rows of $\Omega$.
Now we analyze all possible forms of the truncated rref matrix $E:=\Omega_{:N,:}$ with $R$ columns and $N\leq J\leq R$ rows, for $R\leq 3$:
  • For $R=1$, we must have $E=\begin{bmatrix}x\end{bmatrix}$, so we simply check whether $\mathrm{rank}\left(D_{0}\right)\leq\begin{cases}1&x\neq 0\\ 0&\textrm{else}\end{cases}$.
  • For $R=2$, we either have $E=\begin{bmatrix}1\\ &1\end{bmatrix}$, in which case we check whether $\mathrm{rank}\left(D_{0}\right)\leq 1\ \land\ \mathrm{rank}\left(D_{1}\right)\leq 1$; or $E=\begin{bmatrix}1&x\end{bmatrix},\begin{bmatrix}0&1\end{bmatrix}$, in which case we check whether $\mathrm{rank}\left(D_{0}\right)\leq(\#\textrm{ of nonzeros in }E_{0,:})$.
  • For $R=3$, there are a few more cases, of which one is nontrivial:
    • $E=\begin{bmatrix}1&x&y\end{bmatrix},\begin{bmatrix}0&1&x\end{bmatrix},\begin{bmatrix}0&0&1\end{bmatrix}$: check whether $\mathrm{rank}\left(D_{0}\right)\leq(\#\textrm{ of nonzeros in }E_{0,:})$;
    • $E=\begin{bmatrix}1\\ &1\\ &&1\end{bmatrix}$: check whether $\mathrm{rank}\left(D_{i}\right)\leq 1\ \forall i=0,1,2$;
    • $E=\begin{bmatrix}1&x\\ &&1\end{bmatrix},\begin{bmatrix}0&1\\ &&1\end{bmatrix}$: check whether $\mathrm{rank}\left(D_{i}\right)\leq(\#\textrm{ of nonzeros in }E_{i,:})\ \forall i=0,1$;
    • $E=\begin{bmatrix}1&&x\\ &1&y\end{bmatrix}$ (the nontrivial case):
      If $x=0$ or $y=0$, we can do the same thing as in $\begin{bmatrix}1&x\\ &&1\end{bmatrix}$ or $\begin{bmatrix}0&1\\ &&1\end{bmatrix}$, as the sets of nonzero terms of different rows in $E$ are pairwise disjoint.
      Otherwise, check whether $\exists i\textrm{ s.t. }\mathrm{rank}\left(D_{i}\right)=2$ (notice the $=$ instead of $\leq$ this time):
        if such $i$ exists, enumerate all rank-2 decompositions of $D_{i}=u_{0}\times v_{0}+u_{1}\times v_{1}$, and for each decomposition, check whether $\mathrm{rank}\left(D_{j}-u_{1}\times v_{1}\right)\leq 1$ for the other index $j\neq i$;
        otherwise, $\mathrm{rank}\left(D_{0}\right)\leq 1$ and $\mathrm{rank}\left(D_{1}\right)\leq 1$, so we can force $B_{2,:},C_{2,:}=\overrightarrow{0}$ and solve for $D_{0}=B_{0,:}\times C_{0,:},\ D_{1}=B_{1,:}\times C_{1,:}$;
In contrast to general tensors, rank decomposition of a $n\times n$ matrix (with the decomposition having smallest possible rank) can be computed in $O(n^{3})$ time via Gaussian elimination, so if all relevant checks in this casework pass, it is easy to construct a solution for $B,C$.
It remains to count how many rank-2 decompositions there are for a rank-2 matrix. Fortunately we have such a bound for arbitrary matrix rank which does not have exponential dependence on $n$:
Theorem 3.
Consider an $n\times m$ matrix $M$ over an arbitrary field with rank $r:=\mathrm{rank}\left(M\right)>0$. Let $F_{0}=\mathrm{rref}\left(M\right)_{:r,:}$ and $C_{0}$ satisfy $C_{0}F_{0}=M$. Then for any $n\times r$ matrix $C$ and $r\times m$ matrix $F$ s.t. $CF=M$, there is some invertible $r\times r$ $X$ s.t. $C=C_{0}X^{-1}$ and $F=XF_{0}$.
Proof.
First we note that $C_{0}$ exists and can be constructed via Gaussian elimination by maintaining the composition of all row operations in a separate matrix (and then truncating that matrix to its first $r$ columns). Additionally, $C_{0}$ is unique, because $\mathrm{rank}\left(F_{0}\right)=r$ and $r\leq\min\left(n,m\right)$ force the rows of $F_{0}$ to be linearly independent.
We have $\mathrm{rowspan}\left(M\right)=\mathrm{rowspan}\left(CF\right)\subseteq\mathrm{rowspan}\left(F\right)$ and $\mathrm{rank}\left(F\right)\leq r$, so $\mathrm{rowspan}\left(M\right)=\mathrm{rowspan}\left(F\right)$. Since row reduction and removing all-zero rows preserves row-span, $\mathrm{rowspan}\left(F_{0}\right)=\mathrm{rowspan}\left(\mathrm{rref}\left(F\right)\right)$.
We recall a property of rref matrices. Let $W$ be an arbitrary rref matrix with no all-zero rows. By definition of rref, all terms directly above and below any leading 1 must be 0, so in any linear combination of rows of $W$, leading 1s cannot be canceled out with any other terms. Thus, denoting $l_{i}$ as the column position of the leading 1 at row $i$, the position of the leftmost nonzero element of the row vector $\overrightarrow{v}:=\sum_{i}c_{i}W_{i,:}$ for scalars $c_{i}$ is $l_{k}$ for $k=\min{\left\{i:c_{i}\neq 0\right\}}$. Furthermore, $\overrightarrow{v}_{l_{i}}=c_{i}$, so given an arbitrary linear combination $\overrightarrow{v}$ of the rows of $W$ it is possible to determine the sequence of coefficients $c_{i}$ used to obtain the linear combination. Since the rows of $W$ are linearly independent, there is exactly one such sequence of coefficients $c_{i}$.
Since $\mathrm{rowspan}\left(\mathrm{rref}\left(F\right)\right)$ must contain every row of $F_{0}$, it is possible for a row vector in $\mathrm{rowspan}\left(\mathrm{rref}\left(F\right)\right)$ to have its leftmost nonzero element be at any leading 1’s column in $F_{0}$, forcing $\mathrm{rref}\left(F\right)$ to have its set of column indices of leading 1s be the same as that of $F_{0}$, and thus forcing the leading 1s of $\mathrm{rref}\left(F\right)$ to be in the same set of locations as $\mathrm{rref}\left(F_{0}\right)$. Then, since $\mathrm{rowspan}\left(\mathrm{rref}\left(F\right)\right)$ must contain every row of $F_{0}$, and each row of $F_{0}$ has a 1 at the column index its own leading 1 and a 0 at all column indices of all other leading 1s, $\mathrm{rref}\left(F\right)$ itself must contain each row of $F_{0}$. Because leading 1s in a rref matrix must appear left-to-right as one goes from the top to the bottom row, this is enough to force $\mathrm{rref}\left(F\right)=F_{0}$.
By Gaussian elimination, there is some invertible $r\times r$ matrix $X$ s.t. $F=X\mathrm{rref}\left(F\right)=XF_{0}$, so $M=CF=CXF_{0}$; since the equation $M=AF_{0}$ has exactly one solution for $A$, $CX$ is forced to equal $C_{0}$, so $C=C_{0}X^{-1}$. ∎
The upshot of this result is that over a finite field $\mathbb{F}$, there are exactly $|\mathrm{GL}_{r}\left(\mathbb{F}\right)|=O(|\mathbb{F}|^{r^{2}})$ many rank-$r$ decompositions of a rank-$r$ matrix, and this quantity is independent of the number of rows or columns the matrix has. One can generate all of these decompositions by first constructing $(C_{0},F_{0})$ via Gaussian elimination in $O(n^{3})$ time, then iterate over every $r\times r$ matrix $X$, check whether $X$ is invertible in $O(r^{3})$ time (e.g., by checking if $\det\left(X\right)\neq 0$), then compute $(C,F)=(C_{0}X^{-1},F_{0}X^{-1})$ in $O((n+m)r^{2})$ time. Total running time is then $O(n^{3}+|\mathbb{F}|^{r^{2}}(n+m)r^{5})$. If a list of invertible $r\times r$ matrices and their inverses are precomputed, total running time is $O(n^{3}+|\mathbb{F}|^{r^{2}}(n+m)r^{2})$ instead.
In contrast, the number of rank-$(r+1)$ decompositions of $M$ might already be exponential in the dimensions of $M$ [^3] Consider the case where every element in $M$ is 1..
2.1 Runtime
In our algorithm, we first factor $\overline{T}=M\mathrm{rref}\left(\overline{T}\right)$ and construct $T^{\prime}$ via Gaussian elimination. Since $J=\mathrm{rank}\left(\overline{T}\right)$ must be at most $R$ in order for $T$ to have a rank-$R$ decomposition, we can terminate early if we detect that this condition fails, yielding a running time of $O(Rn^{3})$444 Given a $n\times m$ matrix $A$, we can construct $C,F$ s.t. $A=CF$, $F=\mathrm{rref}\left(A\right)$, and $C$ is invertible, as follows: 1. initialize $(C,F)=(I_{n},A)$ and $r=0$: 2. for $j$ from 00 to $m-1$ inclusive: (a) find some $i\geq r$ s.t. $F_{i,j}\neq 0$ (b) if such $i$ does not exist: skip to the next iteration of the for-loop (c) else: i. swap rows $i$ and $r$ of $F$, and swap columns $i$ and $r$ of $C$ ii. for $k=r+1,\dots n-1$: A. let $s=F_{k,j}$ B. set $F_{k,:}$ to $F_{k,:}-sF_{r,:}$ C. set $C_{:,r}$ to $C_{:,r}+sC_{:,k}$ (here $C_{:,r}$ denotes column $r$ of $C$) iii. increment $r$ The rank of $A$ is exactly the number of times we reach (c). If we only need to return $(C,F)$ under the condition $\mathrm{rank}\left(A\right)\leq R$, for some rank limit $R$, and return a failure statement otherwise, we can terminate early if we ever detect that (c) has been reached more than $R$ times. Thus, this procedure performs $O(Rn(m+n))$ operations. In our main algorithm, $A=\overline{T}$ has shape $n\times n^{2}$, so this procedure runs in $O(Rn^{3})$ time. .
Then we enumerate every possible $U$, of which there are $\leq|\mathbb{F}|^{JR}$ of them. For each $U$, we perform Gaussian elimination on the matrix $\Xi$, which has dimensions $J\times(R+1)$ and has each element in its rightmost column being a $n\times n$ matrix, so we incur $O\left(R^{3}n^{2}\right)$ runtime.
We then ensure there is no $r$ s.t. $\Omega_{r,:}=\overrightarrow{0}$ and $D_{r}\neq O$, which takes $O(J(R+n^{2}))$ time. If this check passes, we proceed to the rref case analysis, during which we compute the ranks of one or more matrices $D_{i}$, each of which takes $O(n^{3})$ time via Gaussian elimination.
If $R=3$ we also have to enumerate all rank-2 decompositions of a rank-2 $n\times n$ matrix $D_{i}$ in the worst case, which takes $O(n^{3}+|\mathbb{F}|^{4}n)$ time (if we precompute all invertible $2\times 2$ matrices over $\mathbb{F}$).
If we pass the case analysis, we construct $B$ and $C$ in $O(n^{3})$ time, then $A:=\left[\sum_{0\leq j<J}M_{i,j}U_{r,j}\right]_{r,i}$ in $O(RJn)$ time, and terminate.
Since we are guaranteed that $J\leq R$ (as otherwise no solution exists), our final runtime is
$O\left(Rn^{3}+\underbrace{|\mathbb{F}|^{R^{2}}}_{\textrm{fix }U}\left(\underbrace{R^{3}n^{2}}_{\textrm{row-reduce }\Xi}+\underbrace{R(n^{2}+R)}_{\textrm{check rows of }\Xi}+Rn^{3}+\left(\begin{cases}n^{3}+|\mathbb{F}|^{4}n&R=3\\ 0&R\leq 2\end{cases}\right)\right)+n^{3}+R^{2}n\right)\\ =O\left(|\mathbb{F}|^{R^{2}}\left(R^{3}n^{2}+Rn^{3}+\left(\begin{cases}n^{3}+|\mathbb{F}|^{4}n&R=3\\ 0&R\leq 2\end{cases}\right)\right)\right)$.
Thus, for any fixed $\mathbb{F}$ and fixed $R$ our algorithm runs in $O(n^{3})$ time, with hidden constant $f(|\mathbb{F}|,R)$ polynomial in $|\mathbb{F}|$ and exponential in $R$; for a loose bound we can set $f(|\mathbb{F}|,R)=|\mathbb{F}|^{R^{2}+4}R^{3}$. For $\mathbb{F}=\mathbb{Z}/2\mathbb{Z}$ and $R=3$, the coefficient on the $n^{3}$ term is $\approx 2^{3^{2}}\cdot(3+1)=4\cdot 2^{9}=2048$.
2.2 Beyond rank 3
We do not know how to make our algorithm run in $\mathrm{poly}(n)$ time for $R=4$. Consider the case $E=\begin{bmatrix}1&&w&x\\ &1&y&z\end{bmatrix}$, $\mathrm{rank}\left(D_{0}\right)=2,\ \mathrm{rank}\left(D_{1}\right)=2$: unlike when $E=\begin{bmatrix}1&&x\\ &1&y\end{bmatrix}$, we cannot simply force $B_{3,:},C_{3,:}=\overrightarrow{0}$, as the system of equations might become too restrictive. One might also notice that $D_{0}+D_{1}=B_{0,:}\times C_{0,:}+B_{1,:}\times C_{1,:}$, so $D_{0}+D_{1}$ has rank $\leq 2$; however, if it happens to have rank 1, then there still might be exponentially many solutions for $(B_{0,:},C_{0,:},B_{1,:},C_{1,:})$, so it is too expensive to simply iterate over all such solutions and check for feasibility of $(B_{2,:},C_{2,:},B_{3,:},C_{3,:})$.
3 With wildcards
When certain elements in $T$ are wildcards (i.e., allowed to have arbitrary value), tensor decomposition becomes much more difficult. In this section we prove Theorem 2, i.e., rank-2 decomposition with wildcards over $\mathbb{Z}/2\mathbb{Z}$ is NP-hard, via a reduction from NAE-3SAT.
Let $\mathcal{I}:=\begin{bmatrix}\begin{bmatrix}1&0\\ 0&0\end{bmatrix}&\begin{bmatrix}0&0\\ 0&1\end{bmatrix}\end{bmatrix}$ be the $2\times 2\times 2$ tensor with 1s along the main diagonal and 0s everywhere else. It can be verified by brute force that the only rank-2 decompositions of $\mathcal{I}$ are $\begin{bmatrix}1\\ 0\end{bmatrix}^{\times 3}+\begin{bmatrix}0\\ 1\end{bmatrix}^{\times 3}$ and $\begin{bmatrix}0\\ 1\end{bmatrix}^{\times 3}+\begin{bmatrix}1\\ 0\end{bmatrix}^{\times 3}$ (even when allowing all-zero vectors). We can parameterize both of these solutions as $\begin{bmatrix}a\\ \neg a\end{bmatrix}^{\times 3}+\begin{bmatrix}\neg a\\ a\end{bmatrix}^{\times 3}$, yielding a variable gadget. By concatenating $n$ copies of $\mathcal{I}$ diagonally [^5] For an example, the result of concatenating $\begin{bmatrix}\begin{bmatrix}a_{0}&a_{1}\\ a_{2}&a_{3}\end{bmatrix}&\begin{bmatrix}a_{4}&a_{5}\\ a_{6}&a_{7}\end{bmatrix}\end{bmatrix}$ and $\begin{bmatrix}\begin{bmatrix}b_{0}&b_{1}\\ b_{2}&b_{3}\end{bmatrix}&\begin{bmatrix}b_{4}&b_{5}\\ b_{6}&b_{7}\end{bmatrix}\end{bmatrix}$ diagonally is $\begin{bmatrix}\begin{bmatrix}a_{0}&a_{1}\\ a_{2}&a_{3}\\ &&&\\ &&&\end{bmatrix}&\begin{bmatrix}a_{4}&a_{5}\\ a_{6}&a_{7}\\ &&&\\ &&&\end{bmatrix}&\begin{bmatrix}&&&\\ &&&\\ &&&\\ &&b_{0}&b_{1}\\ &&b_{2}&b_{3}\end{bmatrix}&\begin{bmatrix}&&&\\ &&&\\ &&&\\ &&b_{4}&b_{5}\\ &&b_{6}&b_{7}\end{bmatrix}\end{bmatrix},$ where the unmarked cells can be filled with arbitrary things. to form the target tensor $T$ and setting all other elements to be wildcards, we force all rank-2 decompositions of $T$ to be of the form [$\ v_{i}\\ \neg v_{i}\\ \vdots\end{bmatrix}^{\times 3}+\begin{bmatrix}\vdots\\ \neg v_{i}\\ v_{i}\\ \vdots\end{bmatrix}^{\times 3}$, for arbitrary variables $v_{i}$.
Since $\mathrm{NAE}\left(a,b,c\right)=\neg(abc\lor(\neg a)(\neg b)(\neg c))$, and the AND-clauses $abc$ and ($(\neg a)(\neg b)(\neg c)$ cannot both be equal at the same time, $\mathrm{NAE}\left(a,b,c\right)=\neg(abc\oplus(\neg a)(\neg b)(\neg c))=(abc+(\neg a)(\neg b)(\neg c)=0)$, where $\oplus$ denotes XOR and is isomorphic to addition mod 2. Because $T$ already forces all elements of vectors in the second outer product to be negations of vectors in the first outer product, we can encode $\mathrm{NAE}\left(v_{i},v_{j},v_{k}\right)$ by setting $T_{2i,2j,2k}=0$; for negative literals, we add 1 to the corresponding tensor index, e.g., $\mathrm{NAE}\left(v_{i},v_{j},\neg v_{k}\right)$ can be encoded by setting $T_{2i,2j,2k+1}=0$.
Thus, we convert a NAE-3SAT instance with $n$ variables and $m$ clauses into a rank-2 decomposition instance on a $2n\times 2n\times 2n$ tensor with $8n+m$ fixed elements and all other elements being wildcards.
Over arbitrary finite fields, there is an issue of scale invariance, i.e., each outer product $a\times b\times c$ can be replaced with $(\alpha a)\times(\beta b)\times(\gamma c)$ for any $\alpha,\beta,\gamma$ s.t. $\alpha\beta\gamma=1$, allowing any one tensor decomposition to be transformed into many other tensor decompositions that evaluate to the same tensor. This scaling freedom makes constructing gadgets for NP-hardness less obvious.
3.1 Rank 1 with wildcards
Here we outline a polynomial-time algorithm for rank-1 decomposition with wildcards over a fixed finite field $\mathbb{F}$.
First, we rewrite the problem as follows: given an arbitrary subset $S\subseteq\left\{0,\dots n-1\right\}^{3}$ and set of elements $T_{i,j,k}\in\mathbb{F}\ \forall(i,j,k)\in S$, find variables $a_{0\dots n-1},b_{0\dots n-1},c_{0\dots n-1}$ s.t. $a_{i}b_{j}c_{k}=T_{i,j,k}\ \forall(i,j,k)\in S$.
When $\mathbb{F}=\mathbb{Z}/2\mathbb{Z}$, $abc=1$ implies $(a,b,c)=(1,1,1)$, so a solution exists if and only if setting
$a_{i}=\lor_{j,k:(i,j,k)\in S}T_{i,j,k},\ b_{j}=\lor_{i,k:(i,j,k)\in S}T_{i,j,k},\ c_{k}=\lor_{i,j:(i,j,k)\in S}T_{i,j,k}$
yields a solution; this is because this construction produces the minimal set of elements in $a\times b\times c$ that are forced to be 1s by the values $T_{i,j,k}$, and this set covers all $(i,j,k)\in S$ that have $T_{i,j,k}=1$. Thus for $\mathbb{F}=\mathbb{Z}/2\mathbb{Z}$ there is an algorithm for rank-1 wildcard decomposition that runs in $O(n^{3})$ time with low constant factor.
For an arbitrary fixed finite field $\mathbb{F}$, we outline a polynomial-time algorithm.
Let $S^{\prime}:=\left\{(i,j,k)\in S:T_{i,j,k}\neq 0\right\}$ denote all points with nonzero values and $I:=\left\{i:\exists(i,j,k)\in S^{\prime}\right\}$, $J:=\left\{j:\exists(i,j,k)\in S^{\prime}\right\}$, $K:=\left\{k:\exists(i,j,k)\in S^{\prime}\right\}$ denote the sets of coordinate values of the nonzero points along each dimension. Since $T_{i,j,k}\neq 0\Rightarrow a_{i},b_{j},c_{k}\neq 0$, we have that $a_{i}\neq 0\ \forall i\in I$, $b_{j}\neq 0\ \forall j\in J$, and $c_{k}\neq 0\ \forall k\in K$, so we check whether $T_{i,j,k}\neq 0\ \forall(i,j,k)\in(I\times J\times K)\cap S$. If this check fails, there is no solution.
Otherwise, we proceed with the observation that all the equations $a_{i}b_{j}c_{k}=T_{i,j,k}$ involve pure products, i.e., expressions of the form $\left(\prod_{i}a_{i}^{p_{i}}\right)\left(\prod_{j}b_{j}^{q_{j}}\right)\left(\prod_{k}c_{k}^{r_{k}}\right)$, for integers $p_{i},q_{j},r_{k}$. If we allow arbitrary multiplication and division between such expressions, all resulting expressions will still be of this form, although some $p_{i},q_{j},r_{k}$ might be negative.
This observation suggests that we can solve this problem somewhat like a linear system of equations: construct a $|S^{\prime}|\times 3n$ integer matrix $E$ and a $|S^{\prime}|\times 1$ $\mathbb{F}$-vector $W$ s.t. each equation $a_{i}b_{j}c_{k}=T_{i,j,k}$ for $(i,j,k)\in S^{\prime}$ is encoded as $\prod_{m}v_{m}^{E_{l,m}}=W_{l}$ for some $l$, where the variables have been re-indexed as $v_{0,\dots n-1}=a_{0,\dots n-1},\ v_{n,\dots 2n-1}=b_{0,\dots n-1},\ v_{2n,\dots 3n-1}=c_{0,\dots n-1}$. Since we restrict ourselves to $S^{\prime}$, each $W_{l}$ is nonzero. During row reductions, elements in $W$ have to be treated differently: when adding between two rows, multiply their values in $W$; when scaling a row by $s$, raise its value in $W$ to the $s$-th power.
However, once we reduce $\left[\begin{array}[]{c|c}E&W\end{array}\right]$ to some normal form, elements in $E$ might not be ±$\pm 1$, so extracting a solution for the variables $v_{i}$ may involve computing $r$-th roots in $\mathbb{F}$. An explicit example of this phenomenon is the tensor
$T=\begin{bmatrix}\begin{bmatrix}1&?\\ ?&1\end{bmatrix},\begin{bmatrix}?&2\\ 1&?\end{bmatrix}\end{bmatrix},$
where ? represents wildcards (equivalently, $S=\left\{(0,0,0),(1,1,0),(0,1,1),(1,0,1)\right\}$, $T_{0,0,0}=1,\ T_{1,1,0}=1,\ T_{0,1,1}=2,\ T_{1,0,1}=1$). Solving this instance is equivalent to solving the equations
$\left\{a_{0}b_{0}c_{0}=1,\ a_{1}b_{1}c_{0}=1,\ a_{0}b_{1}c_{1}=2,\ a_{1}b_{0}c_{1}=1\right\};$
by combining these equations, we get $\frac{(a_{0}b_{1}c_{1})(a_{1}b_{0}c_{1})}{(a_{0}b_{0}c_{0})(a_{1}b_{1}c_{0})}=\frac{2\cdot 1}{1\cdot 1}\Rightarrow\left(\frac{c_{1}}{c_{0}}\right)^{2}=2$, so solving this equation requires us to compute a square root of 2.
To resolve this issue, we use the fact that the multiplicative group of any finite field is cyclic, so if $\mathbb{F}$ is fixed we can find a primitive element (i.e., multiplicative generator) $\mu$ via brute force and define $u_{i},w_{i}$ s.t. $v_{i}=\mu^{u_{i}},\ W_{i}=\mu^{w_{i}}$; then our problem reduces to solving the linear system of equations $Eu=w$ for the vector $u$ over $\mathbb{Z}/(|\mathbb{F}|-1)\mathbb{Z}$.
From [6], solving the matrix-vector equation $xA=b$ over $\mathbb{Z}/N\mathbb{Z}$ for $n\times m$ $A$ can be done in $O(nm^{\omega-1})$ many operations, where $\omega\leq 2.371552$ [8] is the exponent for matrix multiplication, with all integers throughout the algorithm having magnitude $\leq N$. By transposing our system to $u^{\intercal}E^{\intercal}=w^{\intercal}$ and noting that $E$ has dimensions $|S^{\prime}|\times 3n$, we see that solving our system takes $O(n|S^{\prime}|^{\omega-1})$ operations, which is within $O(n^{3\omega-2})$ since $|S^{\prime}|\leq n^{3}$. For fixed $\mathbb{F}$, since all integers used in this algorithm have magnitude $\leq|\mathbb{F}|-1$, each operation takes $O(1)$ time, so our algorithm runs in polynomial time. If we account for the number of wildcards in $T$ in the time complexity, then by letting $k$ be the number of wildcards, $|S^{\prime}|\leq|S|=n^{3}-k$, so we get $O(n(n^{3}-k)^{\omega-1})$.
The main bottleneck in this algorithm is finding a primitive element of $\mathbb{F}$, for which an efficient algorithm is not known [7]. On the other hand, for $\mathbb{F}=\mathbb{F}_{p^{k}}$, the total input size of a problem instance would be linear w.r.t. $\log(p)$ and $k$ [^6] Every element in $\mathbb{F}_{p^{k}}$ is a linear combination of $1,\alpha,\dots,\alpha^{k-1}$ with coefficients in $\mathbb{Z}/p\mathbb{Z}$, for some element $\alpha$, so every element can be specified with a $k$-element vector over $\mathbb{Z}/p\mathbb{Z}$.. Because of this, we are not sure what the computational hardness of rank-1 wildcard decomposition is w.r.t. input size when $p,k$ are not fixed parameters.
3.2 Rank-1 with wildcards for matrices
One may wonder whether matrix rank factorization with wildcards is NP-hard. The issue is that matrix factorizations have an immense degree of freedom, as $CF=M$ implies $(CX^{-1})(XF)=M$, so constructing gadgets seems very difficult even for rank $R=2$.
Additionally, for rank $R=1$, we sketch an $O(n^{3})$ algorithm that does not contain $|\mathbb{F}|$ in its asymptotic running time, in contrast to the analogous problem for 3D tensors. Again we formulate the problem as follows: given a subset $S\subseteq\left\{0,\dots n-1\right\}^{2}$ and set of elements $T_{i,j}\in\mathbb{F}\ \forall(i,j)\in S$, find variables $a_{0,\dots n-1},b_{0,\dots n-1}$ s.t. $a_{i}b_{j}=T_{i,j}$.
Like before, define $S^{\prime}:=\left\{(i,j)\in S:T_{i,j}\neq 0\right\}$, $I:=\left\{i:(i,j)\in S^{\prime}\right\}$, $J:=\left\{j:(i,j)\in S^{\prime}\right\}$. Then $a_{i}\neq 0\ \forall i\in I$ and $b_{j}\neq 0\ \forall j\in J$, so we check whether $T_{i,j}\neq 0\ \forall(i,j)\in(I\times J)\cap S$; if the check fails then there is no solution.
Otherwise, we notice that since every equation $a_{i}b_{j}=T_{i,j}$ constrains the product of two variables, knowing one variable immediately forces the value of the other variable. This motivates us to construct a graph $G$ with vertex set $S^{\prime}$ where there is an edge between every pair of vertices $(i,j),(i^{\prime},j^{\prime})$ s.t. $i=i^{\prime}$ or $j=j^{\prime}$.
Choose an arbitrary $(i_{0},j_{0})\in S^{\prime}$: WLOG we can set $a_{i_{0}}=1,b_{j_{0}}=T_{i_{0},j_{0}}$ because any solution $(a,b)$ to the original problem can be scaled to $(\alpha a,\frac{1}{\alpha}b)$ for any $\alpha\neq 0$. Then do a breadth-first search over $S^{\prime}$ starting from $(i_{0},j_{0})$: when we visit each new vertex $(i,j)\in S$, we will have already determined $a_{i}$ and $b_{j}$, so for each neighbor of the form $(i^{\prime},j)$ we set $a_{i^{\prime}}=T_{i^{\prime},j}/b_{j}$, and for each neighbor of the form $(i,j^{\prime})$ we set $b_{j^{\prime}}=T_{i,j^{\prime}}/a_{i}$. Each neighbor that was not already visited is added to a queue of vertices to visit. Note that because every $T_{i,j}$ we deal with is nonzero, we will never encounter division by 0 in this procedure.
If at any point we update a variable $a_{i}$ or $b_{j}$ that already has a assigned value and set it to a different value, then we have encountered an inconsistency and there is no solution to the original problem. Otherwise, at the end of this procedure we will have searched through the entire connected component of $G$ that contains $(i_{0},j_{0})$, so we do another breadth-first search starting at an arbitrary unvisited vertex in $G$, and repeat until all vertices in $G$ have been visited. If no inconsistencies were ever encountered, we can set all variables $a_{i}$ and $b_{j}$ that were not already assigned a value to 0, and we have a solution.
Since breadth-first search over a graph with $V$ vertices and $E$ edges takes $O(V+E)$ time, and $G$ has $O(n^{2})$ vertices and $O(n^{3})$ edges, our algorithm runs in $O(n^{3})$ time.
The key for correctness is that if two vertices $(i,j),(i^{\prime},j^{\prime})$ are in different connect components, then $(i^{\prime},j^{\prime})$ is not adjacent to any vertex $(x,y)$ in the connected component containing $(i,j)$, so $i^{\prime}\neq x$ and $j^{\prime}\neq y$. This means that $a_{i^{\prime}}$ and $b_{j^{\prime}}$ are not directly or indirectly involved in any equations with elements $T_{x,y}$ for each $(x,y)$ in the connected component of $(i,j)$. Thus, after we finish searching over each connected component, all remaining unassigned variables $a_{i}$ and $b_{j}$ are completely independent of all previously assigned variables, so as we search through the next connected component we will never encounter inconsistencies with variables assigned during a previous connected component. This independence between connected components also allows us to assume WLOG that $a_{i_{0}}=1$ and $b_{j_{0}}=T_{i_{0},j_{0}}$ at each individual connected component.
If we tried generalizing this algorithm to 3D tensors, we would encounter a problem: two vertices would have to match 2 out of 3 coordinates in order to be adjacent to each other, because each equation $a_{i}b_{j}c_{k}=T_{i,j,k}$ constrains the product of three variables instead of two, so two variables have to be known beforehand to force the value of the third one. But then if $(i,j,k)$ and $(i^{\prime},j^{\prime},k^{\prime})$ occupy different components, it is still possible for $(i^{\prime},j^{\prime},k^{\prime})$ to share 1 out of 3 coordinates with a vertex contained in the connected component containing $(i,j,k)$, so different connected components can still share common variables and thus interact with each other in a nontrivial manner.
4 Future directions
We conclude with some open questions:
  • For what $R\geq 4$ can rank-$R$ tensor decomposition (without wildcards) be computed in polynomial time over a fixed finite field?
  • For $R\leq 3$, how low can the constant factor of rank-$R$ decomposition be w.r.t. to the size of the finite field and w.r.t. $R$?
  • What is the parameterized complexity of rank-2 wildcard decomposition w.r.t. the number of wildcard elements in the target tensor? On a related note, how asymptotically low can the number of wildcards in a tensor be while keeping rank-2 wildcard decomposition NP-hard?
  • What is the parameterized complexity of rank-1 wildcard decomposition w.r.t. the size or the characteristic of the underlying finite field?
  • Which of our algorithms be generalized beyond finite fields, such as the integers?
Acknowledgments
We thank Prof. Erik Demaine and TAs Josh Brunner, Lily Chung, and Jenny Diomidova for running MIT 6.5440 Fall 2023, which made this project possible. We also thank them for giving feedback on an earlier version of this report, some of which were requests for certain details and open questions to be addressed, including rank-1 wildcard decomposition.